OpenStudy (anonymous):

I WILL FAN AND MEDAL THE PERSON WHO HELPS ME! Question and options are in the comment

OpenStudy (anonymous):

Simplify completely:\[\frac{ 3x^4+5x^2+2x }{ x }\]

OpenStudy (anonymous):

3x3 + 5x + 2 3x2 + 5x2 + 2x 3x5 + 5x3 + 2x2

OpenStudy (anonymous):

last answer \[\frac{ 3x^3+5x^2+2x }{ x }\]

OpenStudy (mayaal):

the answer is 3x3+5x+2

OpenStudy (queelius):

\[\frac{ 3x^4+5x^2+2x }{ x } = \frac{ 3x^4 }{ x } + \frac{ 5x^2 }{ x } + \frac{ 2x }{ x } = 3x^3+5x+2\]